Help me solve this question asap
(a) A solenoid is of length L and has n turns. i. Write the expression for the magnetic flux density B within the solenoid 11. State, in terms of B, the magnetic flux density at the ends of the solenoid when the current in the solenoid is 1 [2 marks] (b) Two long straight wires are separated by a distance R. Both wires carry the same current / but in opposite direction. Point P is located midway between the two wires. What is the magnetic field strength at P? 5 marks] Two long parallel straight wires X and Y are placed d m apart in a vacuum. The current flowing in wires X and Y are I and 3/respectively and the force per unit length on each wire is kNm1. The current flowing in the wires X and Y are then increased to 2/ and 4/ respectively. What is the distance between the wires so that the force per unit length is 3kN m1? (c) 3 marks
Answer #1

(a) A solenoid is of length 'L' and has 'n' turns.

i. Write an expression for the magnetic flux density 'B' within the solenoid when the current in the solenoid is 'I1'.

Rightarrow The magnetic field within the solenoid has a uniform value 'B' and outside the coils has a value 'zero'.

Using an ampere's law, we have

ointvec{B} . dvec{s} = mu0 Ienc

where, Ienc = current enclosed = I1 (n L)

B L = mu0 I1 n L

B = mu0 n I1

(b) Two long straight wires are separated by a distance R. Both wires carry the same current 'I' but in opposite direction.

Wire 1 Wire 2 (R/2) (R/2

The magnetic field strength at point 'P' which will be given as -

BP = B1 + B2

BP = (mu0 I1 / d1) + (mu0 I2 / d2)

where, d1 = d2 = R/2

I1 = I2 = I

then, we get

BP = [mu0 I / (R/2)] - [mu0 I / (R/2)]

BP = 0

(c) Force between two wires which will be given by -

F / L = mu0 I1 I2 / 2pi d

(k N/m) = (4pi x 10-7 mT/A) (I) (3 I) / 2pi d

(k N/m) = (6 x 10-7 mT) I2 / d

d = (6 x 10-7 mT) I2 / (k N/m)

What is the distance between the wires so that the force per unit length is 3k N/m?

using a formula, we have

F / L = mu0 I1 I2 / 2pi d

(3k N/m) = (4pi x 10-7 mT/A) (2I) (4 I) / 2pi d

(3k N/m) = (16 x 10-7 mT) I2 / d

d = (16 x 10-7 mT) I2 / (3k N/m)